# Policy-Driven World Model Adaptation for Robust Offline Model-based Reinforcement Learning

### v2

@@ 1 Introduction and Related Works
Offline RL [DBLP:journals/corr/abs-2005-01643] leverages offline datasets of transitions, collected by a behavior policy, to train a policy. To avoid overestimation of the expected return for out-of-distribution states, which can mislead policy learning, model-free offline RL methods [DBLP:conf/nips/KumarZTL20, DBLP:conf/nips/AnMKS21] often constrain the learned policy to remain close to the behavior policy. However, acquiring a large volume of demonstrations from a high-quality behavior policy, can be expensive. This challenge has led to the development of offline model-based reinforcement learning (MBRL) approaches, such as [DBLP:conf/nips/YuTYEZLFM20, DBLP:conf/icml/SunZJLYY23, chen2024deep]. These methods train dynamics models from offline data and optimize policies using imaginary rollouts simulated by the models. Notably, the dynamics modeling is independent of the behavior policy, making it possible to learn effective policies from any behavior policy that reasonably covers the state-action spaces.
@@
As detailed in Section 2, offline MBRL typically follows a two-stage framework: first, learning a world model by maximizing the likelihood of transitions in the offline dataset; and second, using the learned model as a surrogate simulator to train RL policies. Unlike in online MBRL [DBLP:conf/icml/RossB12, DBLP:conf/iclr/HafnerLB020, DBLP:conf/nips/ChenAL23], the world model in offline settings is typically not adapted alongside the policy. Moreover, the model training objective, i.e., likelihood maximization, differs from the goal of policy optimization, i.e., maximizing the expected return, leading to the issue of objective mismatch [DBLP:journals/tmlr/Wei0M0C24]. There has been extensive research addressing this mismatch in online MBRL (e.g., [DBLP:conf/nips/Farahmand18, DBLP:conf/nips/GrimmBSS20, DBLP:conf/aaai/NikishinAAB22, DBLP:conf/nips/EysenbachKLS22, DBLP:conf/l4dc/MaSYBJ23, DBLP:conf/icml/Vemula0SBC23]), primarily by directly optimizing the model to increase the current policy’s return in the environment. However, applying these strategies in the offline setting can be problematic. Since the real environment is inaccessible during offline training, optimizing the world model solely to increase returns in imagined rollouts can cause it to diverge from the true dynamics, thereby misleading the policy updates. Comparisons between [DBLP:conf/nips/YangZFZ22] and [DBLP:conf/icml/SunZJLYY23] show that, for offline MBRL, return-driven model adaptation can underperform approaches that do not dynamically update the world model. How to adapt the world model alongside the policy under a unified objective remains an open challenge in offline MBRL. In this paper, we propose such a unified training framework to learn robust RL policies from offline data. Specifically, the objective is formulated as a maximin problem: we maximize the worst-case performance of the policy, where the policy seeks to maximize its expected return while the world model is updated adversarially to minimize it. This direction of model updating is opposite to that used in online MBRL, as conservatism is critical in offline RL [DBLP:journals/corr/abs-2005-01643].
@@
Another motivation behind our algorithm design is that the state-of-the-art model-free and model-base offline RL algorithms lack robustness during deployment, as demonstrated in Figure 1, as the learned policies often overfit to the static dataset or data-driven dynamics models. Changing the objective to maximizing the worst-case performance of the policy can potentially mitigate such issues. While several theoretical studies have investigated robustness in (model-based) RL [DBLP:conf/ijcai/HuangXFZ22, DBLP:conf/nips/PanagantiXKG22, uehara2023pessimisticmodelbasedofflinereinforcement, DBLP:journals/jmlr/ShiC24], the proposed algorithms are not scalable to high-dimensional continuous control tasks. The work most closely related to ours is RAMBO [DBLP:conf/nips/RigterLH22], which also targets robust offline MBRL. However, RAMBO does not formulate the problem as a Stackelberg game, where the policy and world model act as leader and follower in a general-sum game. Also, it does not employ Stackelberg learning dynamics [DBLP:conf/icml/RajeswaranMK20, DBLP:conf/icml/FiezCR20, DBLP:journals/corr/abs-2309-16188, DBLP:journals/corr/abs-2305-17327], which are essential for solving the maximin optimization inherent in robust RL (see Section 2 for details). Additionally, since the world model is updated jointly with the policy, historical rollouts in the replay buffer would become outdated for policy training due to distributional shift. While RAMBO overlooks this issue, we introduce a novel gradient mask mechanism (detailed in Appendix J) to mitigate it while preserving training efficiency.
Our contributions are as follows: (1) We propose a novel offline MBRL algorithm (ROMBRL) that jointly optimizes the model and policy for robustness by solving a constrained maximin optimization problem. We further provide novel theoretical guarantees on the suboptimality gap of the resulting robust policy. (2) To solve this maximin problem, we model it as a Stackelberg game and introduce novel Stackelberg learning dynamics for stochastic gradient updates of the policy and model. To the best of our knowledge, this is the first application of the Stackelberg game framework to offline MBRL. (3) We develop practical implementations of the proposed learning dynamics, including the use of the Woodbury matrix identity for efficient second-order gradient computation and a gradient-mask mechanism that enables off-policy training to improve data efficiency. (4) ROMBRL demonstrates top performance on both the widely-used D4RL MuJoCo benchmark and a challenging Tokamak Control benchmark for nuclear fusion.
@@ 2 Background
Offline Model-based Reinforcement Learning: A Markov Decision Process (MDP) [puterman2014markov] can be described as a tuple $M=\langle\mathcal{S},\mathcal{A},P,R,d_{0},\gamma\rangle$. $\mathcal{S}$ and $\mathcal{A}$ are the state space and action space, respectively. $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ represents the transition dynamics function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{[0,1]}$ defines the reward function, and $d_{0}:\mathcal{S}\rightarrow\Delta_{\mathcal{S}}$ specifies the initial state distribution, where $\Delta_{\mathcal{X}}$ is the set of probability distributions over the space $\mathcal{X}$. $\gamma\in[0,1)$ is a discount factor.
@@
$\max_{\theta}\mathbb{E}_{s\sim d_{0}(\cdot),a\sim\pi_{\theta}(\cdot|s)}\left[Q(s,a;\theta,\phi)\right],\ Q(s,a;\theta,\phi)=\mathbb{E}_{r\sim R_{\phi}(\cdot|s,a),s^{\prime}\sim P_{\phi}(\cdot|s,a)}\left[r+\gamma V(s^{\prime};\theta,\phi)\right]$ (2)
To address uncertainty of the learned world model, offline MBRL methods [DBLP:journals/corr/abs-2410-11234] usually learn an ensemble of world models from $\mathcal{D}_{\mu}$ and apply an ensemble-based reward penalty to $r$ in Eq. (2), discouraging the agent from exploring regions where predictions from the ensemble members exhibit high variance. However, these reward penalty terms are heuristic-based, and as a result, these algorithms lack formal performance guarantees [DBLP:conf/nips/YuKRRLF21].
Learning Dynamics in Stackelberg Games: In a Stackelberg game, the leader (who plays first) and follower aim to solve the following two optimization problems, respectively:
$\min_{x_{1}\in\mathcal{X}_{1}}\{f_{1}(x_{1},x_{2}^{*})\mid x_{2}^{*}\in\arg\min_{x_{2}\in\mathcal{X}_{2}}f_{2}(x_{1},x_{2})\},\ \min_{x_{2}\in\mathcal{X}_{2}}f_{2}(x_{1},x_{2})$ (3)
The authors of [DBLP:conf/icml/FiezCR20] propose update rules for the leader and follower in a class of two-player smooth games defined on continuous and unconstrained $\mathcal{X}_{1},\mathcal{X}_{2}$. Specifically, at each iteration $k$, $x_{1}$ and $x_{2}$ are updated as follows:
$x_{1}^{k+1}=x_{1}^{k}-\eta_{1}^{k}\omega^{k}_{1},\ \omega^{k}_{1}=D_{1}f_{1}(x^{k})-D_{21}f_{2}(x^{k})^{T}(D_{22}f_{2}(x^{k}))^{-1}D_{2}f_{1}(x^{k});$ (4)
$x_{2}^{k+1}=x_{2}^{k}-\eta_{2}^{k}\omega^{k}_{2},\ \omega^{k}_{2}=D_{2}f_{2}(x^{k}).$
@@
The learning target for Stackelberg games is to reach Local Stackelberg Equilibrium (LSE [bacsar1998dynamic]):
Definition 1 (LSE).
@@ 3 Theoretical Results
Our method is based on an robust offline MBRL objective [^1] CPPO-LR, proposed in [uehara2023pessimisticmodelbasedofflinereinforcement], adopts a similar objective function; however, its theoretical analysis – specifically Lemma 6 and Lemma 7 in its Appendix B.2 – is incorrect.: (For simplicity, we merge $P_{\phi}$ and $R_{\phi}$ into a single notation, $P_{\phi}(r,s^{\prime}|s,a)$, in the following discussion.)
@@
The uncertainty set $\Phi=\{\phi\in\mathcal{M}\mid\mathbb{E}_{(s,a)\sim\mathcal{D}_{\mu},(r,s^{\prime})\sim P_{\bar{\phi}}(\cdot|s,a)}\left[\log P_{\bar{\phi}}(r,s^{\prime}|s,a)-\log P_{\phi}(r,s^{\prime}|s,a)\right]\leq\epsilon\}$, where $\bar{\phi}$ is an optimal solution to Eq. (1) (i.e., a maximum likelihood estimator). Intuitively, $\pi_{\theta}$ is trained to maximize its worst-case performance within an uncertainty set of world models, ensuring robust deployment performance.
Define $d_{\theta^{*},\phi^{*}}(s,a)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}d_{\theta^{*},\phi^{*}}^{t}(s,a)$, where $\theta^{*}$ and $\phi^{*}$ represent the parameters of a comparator policy and the true MDP, respectively, and $d_{\theta^{*},\phi^{*}}^{t}(s,a)$ denotes the probability density of the agent reaching $(s,a)$ at time step $t$ when following $\pi_{\theta^{*}}$ in the environment $M_{\phi^{*}}$. Further, as in [uehara2023pessimisticmodelbasedofflinereinforcement], we can define a concentrability coefficient $C_{\phi^{*},\theta^{*}}=\sup_{\phi\in\mathcal{M}}\frac{\mathbb{E}_{(s,a)\sim d_{\theta^{*},\phi^{*}}(\cdot)\left[TV(P_{\phi}(\cdot|s,a),P_{\phi^{*}}(\cdot|s,a))^{2}\right]}}{\mathbb{E}_{(s,a)\sim d_{\mu,\phi^{*}}(\cdot)\left[TV(P_{\phi}(\cdot|s,a),P_{\phi^{*}}(\cdot|s,a))^{2}\right]}}$, where $TV(\cdot,\cdot)$ denotes the total variation distance between two distributions. Then, we have the following theorem [^2] For a discounted finite-horizon MDP with horizon $h$, Theorem 1 remains valid with the following modifications: (1) replacing $\frac{1}{(1-\gamma)^{2}}$ in Eq. (6) with $\frac{(1-\gamma^{h})^{2}}{(1-\gamma)^{2}}$ and (2) replacing $d_{\theta^{*},\phi^{*}}(s,a)$ with $d_{\theta^{*},\phi^{*}}^{h}(s,a)=\frac{1-\gamma}{1-\gamma^{h}}\sum_{t=0}^{h}\gamma^{t}d_{\theta^{*},\phi^{*}}^{t}(s,a)$. The derivation is similar with the one presented in Appendix B.: (Please refer to Appendix B for the proof.)
Theorem 1.
@@
Notably, if $\pi_{\theta^{*}}$ is the optimal policy on $M_{\phi^{*}}$ and $C_{\phi^{*},\theta^{*}}<\infty$, Eq. (6) represents the suboptimality gap of the learned policy $\pi_{\hat{\theta}}$ through Eq. (5). Moreover, Theorem 1 applies to a general function class of $P_{\phi}$ and relies on the assumption that $\phi^{*}\in\Phi$ with probability at least $1-\delta/2$. By specifying the function class of $P_{\phi}$, we can determine the uncertainty range $\epsilon$ to ensure that $\Phi$ includes $\phi^{*}$ with high probability, allowing us to remove this assumption.
Denote $\mathcal{D}^{\mu}_{sa}$ as the number of unique state-action pairs in $\mathcal{D}_{\mu}$ and $N_{sa}$ as the number of transitions in $\mathcal{D}_{\mu}$ that are sampled at $(s,a)$. Further, we define $\widetilde{N}=\max\ \{N_{sa}\mid(s,a)\in\mathcal{D}_{\mu}\}$. For tabular MDPs, the world models (i.e., $P_{\phi}$) follow categorical distributions and we have the following theorem [^3] As defined in [mardia2020concentration], $C_{0}\approx 3.20$ and $C_{1}\approx 2.93$. The expression for $\epsilon$ when $K$ falls into different ranges (e.g., $K\geq N_{sa}C_{0}+2$) can be derived similarly based on Theorem 3 from [mardia2020concentration].: (Please refer to Appendix C for the proof and Appendix E for a discussion on $|\Phi|$.)
Theorem 2.
@@
Further, most recent offline MBRL methods [DBLP:conf/nips/YuTYEZLFM20, DBLP:conf/iclr/LuBPOR22, DBLP:conf/icml/SunZJLYY23] utilize deep neural network-based world models and represent $P_{\phi}(s^{\prime},r\mid s,a)$ as a multivariate Gaussian distribution with a diagonal covariance matrix, $\forall(s,a)$. In this case, we have the following theorem: (Please check Appendix D for a non-asymptotic representation of $\epsilon$ and the proof.)
Theorem 3.

### v3

@@ 1 Introduction and Related Works
Offline RL (Levine et al., 2020) leverages offline datasets of transitions, collected by a behavior policy, to train a policy. To avoid overestimation of the expected return for out-of-distribution states, which can mislead policy learning, model-free offline RL methods (Kumar et al., 2020; An et al., 2021) often constrain the learned policy to remain close to the behavior policy. However, collecting large demonstrations from a high-quality behavior policy, can be expensive. This motivates offline model-based reinforcement learning (MBRL) approaches, such as Yu et al. (2020); Sun et al. (2023); Chen et al. (2024c). These methods train dynamics models from offline data and optimize policies using imaginary rollouts simulated by the models. Notably, the dynamics modeling is independent of the behavior policy, making it possible to learn effective policies from any behavior policy that reasonably covers the state-action spaces.
@@
As detailed in Section 2, offline MBRL typically follows a two-stage framework: first, learning a world model by maximizing transition likelihood in the offline dataset; and second, using the learned model as a surrogate simulator to train RL policies. Unlike in online MBRL (Ross and Bagnell, 2012; Hafner et al., 2020; Chen et al., 2023a), the world model in offline settings is typically not adapted alongside the policy. Moreover, the model training objective (likelihood maximization) differs from policy optimization (maximizing expected return), causing objective mismatch (Wei et al., 2024). There has been extensive research addressing this mismatch in online MBRL (e.g., Farahmand (2018); Grimm et al. (2020); Nikishin et al. (2022); Eysenbach et al. (2022); Ma et al. (2023); Vemula et al. (2023)), primarily by directly optimizing the model to increase the current policy’s return in the environment. However, applying these strategies offline can be problematic. Since the real environment is inaccessible during offline training, optimizing the world model solely to increase returns in imagined rollouts can cause it to diverge from the true dynamics, thereby misleading the policy updates. Comparisons between Yang et al. (2022b) and Sun et al. (2023) show that, for offline MBRL, return-driven model adaptation can underperform approaches that do not dynamically update the world model. How to adapt the world model alongside the policy under a unified objective remains an open challenge in offline MBRL. In this paper, we propose such a unified training framework to learn robust RL policies from offline data. Specifically, the objective is formulated as a maximin problem: we maximize the worst-case performance of the policy, where the policy maximizes its expected return while the world model is updated adversarially to minimize it. This direction of model updating is opposite to online MBRL, as conservatism is critical in offline RL (Levine et al., 2020).
@@
Another motivation behind our algorithm design is that the state-of-the-art model-free and model-based offline RL algorithms lack deployment robustness (Figure 1), as the learned policies often overfit to the static dataset or data-driven dynamics models. Maximizing the worst-case performance of the policy can potentially mitigate such issues. Recent practical robust algorithms like RFQI (Panaganti et al., 2022) and RORL (Yang et al., 2022a) improve robustness via robust value estimation or conservative smoothing, while TRACER (Yang et al., 2024) tackles data corruption through variational inference. However, these methods predominantly operate in a model-free manner, lacking an explicit world model to generalize beyond the dataset support, or focus primarily on training data corruption rather than dynamics mismatch. Closest to our work is RAMBO (Rigter et al., 2022), which also targets robust offline MBRL. However, RAMBO uses alternating updates, implicitly treating the interaction as a symmetric zero-sum game. This approach overlooks the bi-level nature of the problem, where the worst-case world model is an implicit function of the policy. As a result, it lacks theoretical convergence guarantees and often leads to over-conservatism by hedging against all possible dynamics. To address this, we formulate the problem as a Stackelberg game, capturing the asymmetric leader-follower relationship. We employ Stackelberg learning dynamics (Rajeswaran et al., 2020; Fiez et al., 2020; Zhou and Qu, 2023; Chen et al., 2023c), which are essential for properly solving the inherent maximin optimization with convergence guarantees (see Section 2 for details). Additionally, since the world model is updated jointly with the policy, historical rollouts in the replay buffer would become outdated for policy training due to distributional shift. While RAMBO overlooks this issue, we introduce a novel gradient mask mechanism (detailed in Appendix J) to mitigate it while preserving training efficiency.
Our contributions are: (1) We propose ROMBRL, a novel offline MBRL algorithm that jointly optimizes the model and policy for robustness via constrained maximin optimization, providing novel theoretical guarantees on the policy’s suboptimality gap. (2) We model this problem as a Stackelberg game and introduce novel Stackelberg learning dynamics for stochastic gradient updates. This is the first application of the Stackelberg framework to offline MBRL. (3) We develop practical implementations, including using the Woodbury matrix identity for efficient second-order gradients and a gradient-mask mechanism for data-efficient off-policy training. (4) ROMBRL achieves state-of-the-art performance and robustness on the D4RL MuJoCo benchmark under noise and the challenging Tokamak Control benchmark. In clean settings, it performs on par with SOTA algorithms and outperforms RAMBO, a robust baseline.
@@ 2 Background
Offline Model-based Reinforcement Learning: A Markov Decision Process (MDP) (Puterman, 2014) can be described as a tuple $M=\langle\mathcal{S},\mathcal{A},P,R,d_{0},\gamma\rangle$. $\mathcal{S}$ and $\mathcal{A}$ are the state and action space. $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ represents the transition dynamics function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{[0,1]}$ defines the reward function, and $d_{0}:\mathcal{S}\rightarrow\Delta_{\mathcal{S}}$ specifies the initial state distribution, where $\Delta_{\mathcal{X}}$ is the set of probability distributions over the space $\mathcal{X}$. $\gamma\in[0,1)$ is a discount factor.
@@
$\begin{split}&\max_{\theta}\mathbb{E}_{s\sim d_{0}(\cdot),a\sim\pi_{\theta}(\cdot|s)}\left[Q(s,a;\theta,\phi)\right],\\ &Q(s,a;\theta,\phi)=\mathbb{E}_{\begin{subarray}{c}r\sim R_{\phi}(\cdot|s,a)\\ s^{\prime}\sim P_{\phi}(\cdot|s,a)\end{subarray}}\left[r+\gamma V(s^{\prime};\theta,\phi)\right]\end{split}$ (2)
To address uncertainty of the learned world model, offline MBRL methods (Chen et al., 2024a) usually learn an ensemble of world models from $\mathcal{D}_{\mu}$ and apply an ensemble-based reward penalty to $r$ in Eq. (2), discouraging the agent from exploring regions where ensemble predictions exhibit high variance. However, these reward penalty terms are heuristic-based, and as a result, these algorithms lack formal performance guarantees (Yu et al., 2021).
Learning Dynamics in Stackelberg Games: In a Stackelberg game, the leader (who plays first) and follower aim to solve the following optimization problems, respectively:
$\begin{split}&\min_{x_{1}\in\mathcal{X}_{1}}\left\{f_{1}(x_{1},x_{2}^{*})\;\middle|\;\begin{aligned} x_{2}^{*}\in\arg\min_{x_{2}\in\mathcal{X}_{2}}f_{2}(x_{1},x_{2})\end{aligned}\right\},\\ &\min_{x_{2}\in\mathcal{X}_{2}}f_{2}(x_{1},x_{2})\end{split}$ (3)
Fiez et al. (2020) propose update rules for the leader and follower in a class of two-player smooth games defined on continuous and unconstrained $\mathcal{X}_{1},\mathcal{X}_{2}$. Specifically, at each iteration $k$, $x_{1}$ and $x_{2}$ are updated as follows:
$x_{1}^{k+1}$ $=x_{1}^{k}-\eta_{1}^{k}\omega^{k}_{1},$ (4)
$\omega^{k}_{1}$ $=D_{1}f_{1}(x^{k})$
$\quad-D_{21}f_{2}(x^{k})^{T}(D_{22}f_{2}(x^{k}))^{-1}D_{2}f_{1}(x^{k});$
$x_{2}^{k+1}$ $=x_{2}^{k}-\eta_{2}^{k}\omega^{k}_{2},\quad\omega^{k}_{2}=D_{2}f_{2}(x^{k}).$
@@
The learning target for Stackelberg games is to reach Local Stackelberg Equilibrium (LSE (Başar and Olsder, 1998)):
Definition 2.1 (LSE).
@@ 3 Theoretical Results
Our method is based on an robust offline MBRL objective [^1] CPPO-LR, proposed in Uehara and Sun (2023), adopts a similar objective function; however, its theoretical analysis – specifically Lemma 6 and Lemma 7 in its Appendix B.2 – is incorrect.: (For simplicity, we merge $P_{\phi}$ and $R_{\phi}$ into a single notation, $P_{\phi}(r,s^{\prime}|s,a)$, in the following discussion.)
@@
The uncertainty set $\Phi$ is defined as: $\Phi=\{\phi\in\mathcal{M}\mid\mathbb{E}_{(s,a)\sim\mathcal{D}_{\mu}}[D_{\mathrm{KL}}(P_{\bar{\phi}}(\cdot|s,a)\,\|\,P_{\phi}(\cdot|s,a))]\leq\epsilon\}$, where $\bar{\phi}$ is an optimal solution to Eq. (1) (i.e., a maximum likelihood estimator). Intuitively, $\pi_{\theta}$ is trained to maximize its worst-case performance within an uncertainty set of world models, ensuring robust deployment performance.
Define $d_{\theta^{*},\phi^{*}}(s,a)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}d_{\theta^{*},\phi^{*}}^{t}(s,a)$, where $\theta^{*}$ and $\phi^{*}$ represent the parameters of a comparator policy and the true MDP, respectively, and $d_{\theta^{*},\phi^{*}}^{t}(s,a)$ denotes the probability density of the agent reaching $(s,a)$ at time step $t$ when following $\pi_{\theta^{*}}$ in the environment $M_{\phi^{*}}$. Further, as in Uehara and Sun (2023), we can define a concentrability coefficient $C_{\phi^{*},\theta^{*}}=\sup_{\phi\in\mathcal{M}}\frac{\mathbb{E}_{(s,a)\sim d_{\theta^{*},\phi^{*}}(\cdot)\left[TV(P_{\phi}(\cdot|s,a),P_{\phi^{*}}(\cdot|s,a))^{2}\right]}}{\mathbb{E}_{(s,a)\sim d_{\mu,\phi^{*}}(\cdot)\left[TV(P_{\phi}(\cdot|s,a),P_{\phi^{*}}(\cdot|s,a))^{2}\right]}}$, where $TV(\cdot,\cdot)$ denotes the total variation distance between two distributions. Then, we have the following theorem [^2] For a discounted finite-horizon MDP with horizon $h$, Theorem 3.1 remains valid with the following modifications: (1) replacing $\frac{1}{(1-\gamma)^{2}}$ in Eq. (6) with $\frac{(1-\gamma^{h})^{2}}{(1-\gamma)^{2}}$ and (2) replacing $d_{\theta^{*},\phi^{*}}(s,a)$ with $d_{\theta^{*},\phi^{*}}^{h}(s,a)=\frac{1-\gamma}{1-\gamma^{h}}\sum_{t=0}^{h}\gamma^{t}d_{\theta^{*},\phi^{*}}^{t}(s,a)$. The derivation is similar with the one presented in Appendix B.: (Please refer to Appendix B for the proof.)
Theorem 3.1.
@@
$\begin{split}&J(\theta^{*},\phi^{*})-J(\hat{\theta},\phi^{*})\leq\frac{\sqrt{C_{\phi^{*},\theta^{*}}}}{(1-\gamma)^{2}}\times\\ &\quad\sqrt{4\epsilon+c\left(\sqrt{\frac{\log(2|\Phi|/\delta)}{N}}+\frac{\log(2|\Phi|/\delta)}{N}\right)},\end{split}$ (6)
@@
Notably, if $\pi_{\theta^{*}}$ is the optimal policy on $M_{\phi^{*}}$ and $C_{\phi^{*},\theta^{*}}<\infty$, Eq. (6) represents the suboptimality gap of the learned policy $\pi_{\hat{\theta}}$ through Eq. (5). Moreover, Theorem 3.1 applies to a general function class of $P_{\phi}$ and relies on the assumption that $\phi^{*}\in\Phi$ with probability at least $1-\delta/2$. By specifying the function class of $P_{\phi}$, we can determine the uncertainty range $\epsilon$ to ensure that $\Phi$ includes $\phi^{*}$ with high probability, allowing us to remove this assumption.
Denote $\mathcal{D}^{\mu}_{sa}$ as the number of unique state-action pairs in $\mathcal{D}_{\mu}$ and $N_{sa}$ as the number of transitions in $\mathcal{D}_{\mu}$ that are sampled at $(s,a)$. Further, we define $\widetilde{N}=\max\ \{N_{sa}\mid(s,a)\in\mathcal{D}_{\mu}\}$. For tabular MDPs, the world models (i.e., $P_{\phi}$) follow categorical distributions and we have the following theorem [^3] As defined in Mardia et al. (2020), $C_{0}\approx 3.20$ and $C_{1}\approx 2.93$. The expression for $\epsilon$ when $K$ falls into different ranges (e.g., $K\geq N_{sa}C_{0}+2$) can be derived similarly based on Theorem 3 from Mardia et al. (2020).: (Please refer to Appendix C for the proof and Appendix E for a discussion on $|\Phi|$.)
Theorem 3.2.
@@
Further, most recent offline MBRL methods (Yu et al., 2020; Lu et al., 2022; Sun et al., 2023) utilize deep neural network-based world models and represent $P_{\phi}(s^{\prime},r\mid s,a)$ as a multivariate Gaussian distribution with a diagonal covariance matrix, $\forall(s,a)$. In this case, we have the following theorem: (Please check Appendix D for a non-asymptotic representation of $\epsilon$ and the proof.)
Theorem 3.3.